OpenStudy (anonymous):

Solve the following:

OpenStudy (anonymous):

\[2^{x+3}=7\]

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

def: \[\ln^x(a) = x \ln (a)\] take the natural logs of both sides to get \[\ln^(x+3)(2) = \ln(7)\]having problems with my equation editor...i'll try to be as clear as possible. the result should be \[(x+3) \ln(2) = \ln (7)\] by using the definition above. then divide a \[\ln (2)\] on both sides to get \[(x+3) = (\ln (7)/\ln (2))\] isolate x by subtracting a 3 from both sides to get \[x = (\ln(7)/\ln(2)) - 3\] use a calculator to get your results